OpenStudy (anonymous):

how can i solve it this differential equation xcos^2ydx-ycos^2xdy=0

OpenStudy (anonymous):

separate variables .. i dont think you can get explicit solution for this .. separating variables \[\Large \frac{y}{\cos^2(y)}*dy=\frac{x}{\cos^2(x)}*dx\] or can be written as \[\Large \sec^2(y)ydy=\sec^2(x)xdx\] just integrate both sides by applying by parts integration.

OpenStudy (anonymous):

when i use by parts integration formula then answer doesn't come